The difference of two integers is 9. Five times the smaller is 7 more than 3 times the larger. Find the numbers.
I got as far as writing the equations:
x - y = 9
3x + 7 = 5y

Answer by advanced_Learner(501)

3x+7=5x-45
52=2x
x=26



x-y=9
26-y=9
17=y